OpenStudy (jiteshmeghwal9):

solve the equation:- \[x=3y\]\[\frac{ x }{ 3 }+y=34\]

OpenStudy (jiteshmeghwal9):

@ganeshie8 plz help :)

OpenStudy (anonymous):

x=66 y=22

OpenStudy (jiteshmeghwal9):

i need help in solving

OpenStudy (jiteshmeghwal9):

ans.-x=51,y=17

OpenStudy (jiteshmeghwal9):

these r the answers given in book

OpenStudy (jiteshmeghwal9):

x-3y=0 (x/3)+y=34

OpenStudy (jiteshmeghwal9):

x-3y=0 x+3y=102 ohh! i gt it :)

OpenStudy (rational):

substitute x = 3y in second equation

OpenStudy (anonymous):

i am so sorry the ans is x=51 y=17

OpenStudy (jiteshmeghwal9):

2x=102 x=51 &y=17 :)

OpenStudy (jiteshmeghwal9):

thx for ur help @esamalaa @rational :)

OpenStudy (rational):

np :)

OpenStudy (anonymous):

anytime